### v1

# Enhancing Conversational Agents from Open-Source Large Language Models with Illocutionary Force and Document-Based Knowledge Retrieval

Abstract
In this paper, we first present a novel way of computationally analysing and extracting illocutionary forces from dialogue using Bert-based Large Language Models, and demonstrate how these features impact the response of a conversational agent guided by a document-based knowledge bank demonstrated by a bespoke web conversational chat agent system developed. Our proposed illocutionary force extraction and classification technique is the first of its kind using the Argument Interchange Format (AIF) Dataset, showing an improved performance compared to two methods for carrying out similar tasks with a macro F1 of approximately 45%. When we evaluated the system based on 2 knowledge files, with 2 user queries each, across 5 open-source large language models (LLMs) using 10 standard metrics we found out that larger open-source models, such as Llama2:13b and Llama3-chatqa-latest, demonstrated an improved alignment when the user illocutionary force was included with their query, achieving higher QA and linguistic similarity scores. The smaller models on the other hand like Tinyllama:latest showed an increased perplexity and mixed performance, which explicitly indicated struggles in processing queries that explicitly included illocutionary forces. The results from the analysis highlight the potential of illocutionary force to enhance conversational depth while underscoring the need for model-specific optimizations to address increased computational costs and response times.
1 Introduction
The spontaneous evolution and transformation of conversational agents, ignited by the improvements in LLMs has transformed the way in which humans interact and respond with technology. Due to factors like adaptability, accessibility, and performance, open-source LLMs have gained significant attention. Regardless of these strengths, these open-source models often struggle with refined linguistic features such as illocutionary force which forms the the intended speech act behind a statement or locution, creating a gap in better understanding the speaker or user’s communicative intent. Handling this limitation is essential for improving the relevance, coherence, and responsiveness of conversational agents.
This research work introduces a novel framework for analyzing and extracting illocutionary forces behind speakers’ or authors’ locutions using BERT-based transformer architectures and integrating these extracted features into a custom document-based conversational system. When we combined the AIF dataset with our novel dynamic class weight balance technique to tackle class imbalance, we achieved a better macro F1 score which outperformed existing methods by approximately 10% to 15%. Additionally, this research work explores the impact of integrating the user illocutionary force into the user queries for the conversational agents which leverages document-based knowledge retrieval to improve the quality of the response and effectively evaluate its performance using standard metrics which cross-reference the chat agent response to a pre-existing answer stored somewhere (information from a custom file). The experiments we conducted showed varying performance across the 5 open-source models used across the 11 standard evaluation metrics used. The impacts of this research are twofold:
  1. 1.
    It exposes the normally overlooked role and impact of illocutionary force in discourse analysis and its broad impact on conversational agents.
  2. 2.
    It illustrates the potential of combining open-source LLMs with a custom-based document retrieval framework to improve Human-Computer Interaction (HCI).
This work paves the way for more intuitive, context-aware, and adaptive conversational systems.
2 Literature Review
It has been seen that open-source LLMs such as LLaMA [17], Mistral [18], and TinyLLama among others have risen as a relevant and needful alternative to closed-source models because of the scalability, flexibility, and creativity in which they offer. These open-source models have demonstrated imminent relevance in various NLP and machine learning tasks such as natural language generation (NLG) and natural language understanding (NLU). While praising these models for their remarkable ability to capture semantic and syntactic nuances in various tasks, they still struggle with capturing pragmatic nuances such as illocutionary force [6]. Addressing this limitation opens up a less trivial opportunity for additional linguistic and contextual analysis interfaces to improve conversational applications.
To give a background Illocutionary Force in Text Analysis, Illocutionary force is often referred to as the communicative intent behind a statement or locution, as defined in speech act theory [1] [15]. Thus far, there has been a lot of visible significance of the concept of illocutionary forces theoretically, but very limited practical application to NLP. It is seen that more focused text analysis for sentiment, emotion, and intent classification has been enabled by recent improvements in transformer-based architecture models, such as BERT and its derivatives [5]. However, illocutionary force classification remains underexplored and overlooked in computational studies. The core of this study is to build on existing foundations of the speech act theory, leveraging the AIF dataset to model illocutionary forces in dialogue structures, thereby filling in a critical research gap in text analysis for conversational systems.
3 Illocutionary Force Modelling
3.1 Dataset
To give a summary of the data structure and components of the data set worked with, we made use of various annotated corpus stored in the Argument Interchange Format Database (AIFdb) [8] with a JSON format. The key data components in the structure are the nodes which represent various discourse-argumentative units, edges which show how various nodes are inter-connected, and the locution nodes which give details about the participants in the discourse. The Corpora used were US2016, QT30, MM123, and MM2012. We represent these four corpora with ‘UQMM’. When analyzing the distribution of the illocutionary forces we first had all these class lists:
Original_Labels = [
    ’Asserting’, ’Disagreeing’, ’Pure Questioning’, ’Agreeing’,
    ’Assertive Questioning’, ’Rhetorical Questioning’, ’Popular’,
    ’Default Illocuting’, ’Arguing’, ’Assertive Challenging’,
    ’Challenging’, ’Ironic Asserting’, ’Rhetorical Challenging’,
    ’Offering’, ’Pure Challenging’, ’Restating’, ’Ad verecundiam’,
    ’Directive Questioning’, ’Order’, ’Conceding’, ’YA’, ’Expressing’,
    ’Analysing’, ’Conceding’
Following a Qualitative analysis, we pruned these labels [“YA”, “Expressing”, “Offering”, “Ad verecundiam”, “Order”, “Analysing”, “Conceding”, “Popular Conceding”], as some were errors from annotation and others were not needed.
At the final analysis stage, we noticed that some class labels were closely related, we then decided to merge closely related classes together to minimize bias during machine learning. These were “Directive Questioning into Pure Questioning”, “Ironic Asserting into Asserting”, and “Assertive Challenging, Pure Challenging, Rhetorical Challenging into Challenging”. We show three lists below which consist of the class lists after the running and merging pre-processing task.
Original_Labels = [
    ’Asserting’, ’Disagreeing’, ’Pure Questioning’, ’Agreeing’,
    ’Assertive Questioning’, ’Rhetorical Questioning’, ’Popular’,
    ’Default Illocuting’, ’Arguing’, ’Assertive Challenging’,
    ’Challenging’, ’Ironic Asserting’, ’Rhetorical Challenging’,
    ’Offering’, ’Pure Challenging’, ’Restating’, ’Ad verecundiam’,
    ’Directive Questioning’, ’Order’, ’Conceding’, ’YA’, ’Expressing’,
    ’Analysing’, ’Conceding’
Prunning_Results = [
    ’Asserting’, ’Disagreeing’, ’Pure Questioning’, ’Agreeing’,
    ’Assertive Questioning’, ’Rhetorical Questioning’, ’Conceding’,
    ’Default Illocuting’, ’Arguing’, ’Assertive Challenging’,
    ’Challenging’, ’Ironic Asserting’, ’Rhetorical Challenging’,
    ’Challenging’, ’Pure Challenging’, ’Restating’,
    ’Directive Questioning’, ’Pure Questioning’
Merged_Labels = [
    ’Asserting’, ’Disagreeing’, ’Pure Questioning’, ’Agreeing’,
    ’Assertive Questioning’, ’Rhetorical Questioning’,
    ’Default Illocuting’, ’Arguing’, ’Challenging’,
    ’Challenging’, ’Asserting’, ’Challenging’,
    ’Challenging’, ’Restating’
Table 2 below further shows the class distributions across the corpus combinations experimented with.
Prunned Labels UQ UQMM Merged Labels UQMM
Agreeing 108 110 Agreeing 110
Arguing 33 33 Arguing 33
Asserting 13820 27449 Asserting 27476
Assertive Challenging 38 45 Assertive Questioning 629
Assertive Questioning 493 629 Challenging 118
Challenging 42 42 Default Illocuting 68
Conceding 2 4 Disagreeing 26
Default Illocuting 66 68 Pure Questioning 1471
Disagreeing 26 26 Restating 10
Ironic Asserting 26 27 Rhetorical Questioning 404
Popular Conceding 2 30
Pure Challenging 10 15
Pure Questioning 1382 1471
Restating 5 10
Rhetorical Challenging 3 16
Rhetorical Questioning 313 404
Table 1: Class Counts
3.2 Pre-processing
Our preprocessing stage was divided into four major stages seen below:
  • Corpus files concatenation: Each file constituent of the corpus stored in the AIFdb was merged into one large file consisting of all the node elements identifiers (nodeID, text, type, timestamp) and unique values in one list with the key denoted as “nodes”, and all the edges elements identifiers (edgeID, fromID, toID) and values in another list with the key denoted as “edges”.
  • Features extraction: At this stage, we attempt to extract all the features (Locutions and Illocutionary forces) by looping through the data to identify all the locutions nodes denoted by ‘L’ which is a fromID and the toID is an illocutionary force node denoted by ‘YA’. We defined the fromID key in the edges list as ‘from_id’, and the toID key as ‘to_id’. Finally, we iterate through all the edges, and for each fromID which had a type value of L in the nodes list, and the toID of the same edge map with a type value of YA in the nodes list, we add the L text value, and the YA text value to an initialized list of Locutions and Illocutionary Forces while ensuring that their indexes match. This iteration continues until all the conditions are met through the data.
  • Speaker name stripping: After deriving all the locutions and illocutionary forces linked together, at the stage, we progressed to using a regex match to strip out the Speaker names from the locutions which could introduce noise during machine learning.
  • Data restructuring: We finally generated a resultant file which was the training data to be used. These were the locutions and illocutionary forces from the initialized lists with their from and to ID values appended into various JSON maps with a list denoted by illocutions.
    Figure 1: Training Data Structure
3.3 Model Fine-tuning
Using an inverse frequency fine-tuning novel approach of transformer models for handling class imbalance [11], we fine-tuned a pre-trained Bert-base, Distilbert, Roberta-base, and Roberta-large model on an 80:20 train-test split with learning_rate=2e-5, epoch=20, batch_size=50, max_length=50, seed=42, and using cross-entropy loss function. The ‘L_text’ key value consisting of the author utterance from the result of the preprocessing served as the model input while the ‘YA_text’ key value consisting of the illocutionary forces served as the class labels.
Corpus Model Precision Recall F1 Score Accuracy
UQMM Fine-tuned Bert-base 0.43 0.37 0.39 0.64
Fine-tuned Distilbert 0.39 0.34 0.36 0.63
Fine-tuned RoBERTa-base 0.59 0.48 0.52 0.65
Fine-tuned RoBERTa-large 0.57 0.52 0.54 0.66
Table 2: Performance across Transformer Models on Standard Approaches
3.4 Comparison and Evaluation
Comparing our fine-tuning modelling approach with other methods, we further explored two other methods which have been used to address an identical task to ours. These are:
  • Extreme multi-label text classification using CNN: This approach was used for classifying multiple labels from very large label classes to various documents using Convolutional Neural Networks (CNN) forming the core of their work on Extreme multi-label text classification (XMTC) using CNN [10].
  • Few-Shot Using Transformers with Low-Rank Adaptation: Very recently, Jairam et al. [12] proposed an approach which incorporated transformer models fine-tuned and refined with Low-Rank Adaptation of Large Language Models (LoRA) and Quantized Low-Rank Adaptation of Large Language Models (QLoRA) with few-shot learning on six Indian English accents in the Indic Accent imbalanced Dataset with distribution of (Tamil at 1,640, Malayalam at 1,563, 175 Telugu at 1,614, Gujarati at 298, Hindi at 827, Kannada at 1,486) which showed an Accuracy of up to 94% with trainable parameters set at 5%. We only compared with LoRA for this research.
The table below shows the performance of our approach as compared to the two approaches above.
Method Model Epoch Precision Recall F1 Score Accuracy
Inverse Class Frequency Fine-tuned Bert-base 20 0.43 0.37 0.39 0.64
Fine-tuned Distilbert 20 0.39 0.34 0.36 0.63
Fine-tuned RoBERTa-base 20 0.59 0.48 0.52 0.65
Fine-tuned RoBERTa-large 20 0.57 0.52 0.54 0.66
Other Methods XMTC CNN 10 0.25 0.26 0.25 0.50
BERT + LoRA 10 0.04 0.01 0.06 0.38
BERT + LoRA 20 0.08 0.05 0.16 0.44
BERT + LoRA 40 0.18 0.21 0.19 0.63
RoBERTa + LoRA 10 0.12 0.14 0.12 0.47
RoBERTa + LoRA 20 0.20 0.21 0.19 0.57
RoBERTa + LoRA 40 0.35 0.36 0.35 0.69
Table 3: Performance across methods
4 System Overview
Figure 2: Bespoke Conversational System
The Web Application in Figure 2 was custom-developed for the purpose of this research. It offers the public the ability to carry out research and experiments with illocutionary forces, knowledge base, and open-source LLMs. We give an overview of three key features used for this research below:
4.1 Model Selection
This gives end users the ability to choose from a range of various open-source LLM. In the case of this research, we only explored 5 models (llama2:13b Tinyllama:latest Llama3-chatqa-latest Llama3-latest Mistral:latest)
4.2 Illocutionary Force Identification
This component forms the backbone of this research as it gives the ability to either include the user illocutionary force from their query alongside their query to the conversational agent or just pass their query alone to the conversational agent to generate a response.
4.3 Document-Based Knowledge Retrieval
In this integration, we aimed to incorporate illocutionary force into response generation from information during document retrieval for enhancing contextual alignment between the retrieved information and the user intent. The end users have the ability to upload and select various knowledge files in either .txt or .pdf formats to use as the conversational agent’s knowledge to guide its response.
5 Experiment Design
To fully understand and expose the impacts of including the illocutionary force of a user query alongside their query, we set up two key experiments.
Experiment 1: In the first experiment, we used two custom documents seen in section 5.1 below as a knowledge base of the model to restrict its knowledge to only the information in the document. We provided two queries for each of the 5 models while assessing without including the users’ illocutionary force, and evaluated the performance of each model based on 10 metrics seen in section 5.5.
User QueryConversational AgentOllama ModelCustom DocumentGenerated Response
The User Query serves as the input to the conversational AI system. This conversational AI system is driven by 5 Ollama Model which is guided by Custom Documents to produce the required Generated Response based on the prompt designed.
Experiment 2: In the second experiment, we used two custom documents seen in section 5.1 below as a knowledge base of the model to restrict its knowledge to only the information in the document. We provided two queries and the illocutionary forces extracted from the queries for each of the 5 models and evaluated the performance of each model based on 10 metrics seen in section 5.5.
User Query + User Illocutionary ForceConversational AgentOllama ModelCustom DocumentGenerated Response
The User Query and Illocutionary Force serves as the input to the conversational AI system. This conversational AI system is driven by 5 Ollama Model which is guided by Custom Documents to produce the required Generated Response based on the prompt designed.
5.1 Chatbot Knowledge Base Data
5.2 Prompt Design
In designing the prompt for the conversational agent, four parameters were used. The ‘user_input’ takes the user query argument from the front-end; The ‘context’ consists of a list of the previous conversation between the user and the conversational agent with indicators of users and assistants so that the agent can understand which previous conversation was the user and which was from itself just to ensure coherence; The ‘knowledge_text’ which consist of the text content of the knowledge document selected from the front-end, and the ‘illocutionary_force’ which consist of the illocutionary force class assigned to the user query after inference from the model stored on hugging-face. The inclusion of the ‘illocutiomnary_force’ is been controlled from the front-end hence the condition in the prompt which is initialised to ‘True’ but updates based on the choice of the end-user. We show these representations below in a pseudo-code.
else:
# Formulate the updated instruction
instruction = f"""
Respond to the user’s query: "{user_input}" while considering the relevant context from previous conversations.
Please focus on providing a response based on the latest exchange, without repeating the entire conversation history.
If needed, use the context provided below for reference.
Relevant context: "{context}"
Use only the information provided in this document to respond: ’{knowledge_text}’.
if include_illocutionary_force:
instruction += f"Consider the user’s communicative intent while responding, characterized by the illocutionary force: ’{illocutionary_force}’."
# Call Ollama local model to generate response
response = ollama.generate(
model=selected_model,
prompt=instruction,
options={
"temperature": 1, # No randomness
"top_p": 1, # Consider all possible tokens
"top_k": 1, # Choose the highest-probability token
"num_predict": 300, # Set a fixed length for output
"seed": 42, # Ensure reproducibility
"num_ctx": 4096, # Maximize context window for tracking history
"repeat_last_n": -1,
"repeat_penalty": 1.5,
"mirostat_tau": 1.0
stream=False, # Disable streaming for consistent output
raw=False # Enable prompt formatting for structured input
response_content = response.get(’response’, ’’).strip()
5.3 User Queries for each chatbot knowledge base
For Child Health Document (001):
USER QUERY 1: What can I expect from my young child’s development?
USER QUERY 2: That’s great, thanks for helping.
For Poverty in a rising Africa Document (004):
USER QUERY 1: It is quite sad to see the poverty level in Africa. What do you think can be done to solve it?
USER QUERY 2: Brilliant, that sounds amazing.
5.4 Conversational agent response evaluation metrics
For Relevance and Quality:
  • Recall-Oriented Understudy for Gisting Evaluation (ROUGE-1): Rouge is a conversation metric which is used to evaluate the quality of a response (in most cases - a chatbot) comparing its overlapping unigram with a standardise reference response (normally a human-generated response) [9]. In this work, ROUGE-1 is used to measure the presence of certain keywords in the generated response of the chatbot with reference to the custom document used to guide its response. A high ROUGE-1 score indicates that the chatbot’s response captures the necessary relevant information from the reference document.
  • ROUGE-2: Unlike ROUGE-1 which measures unigram with regards to the reference source document, ROUGUE-2 measures the bigram to ensure coherence and contextual alignment between the chatbot response regarding the custom document [16]. We can also note that it captures both syntactic and contextual relationships making it ideal for evaluating phrase-level accuracy.
  • ROUGE-L: ROUGE-L on the other hand evaluates the longest common subsequence between the generated chatbot response and the custom document been used to guide its response. It covers a sentence level, unlike ROUGE-1 and ROUGE-2, making it a good evaluator for syntactic and discourse-level performance. Lin et al., in 2004 [9] argued that ROUGE-L was better for capturing sentence level coherence as compared with ROUGE-1 and ROUGE-2. A higher ROUGE-L score indicates that the chatbot response maintains a similar structure as compared to the reference custom document.
  • Question-Answer Reference (QA-Ref): This measure is used for typical Questions-Answers scenarios where the answer given by the chatbot agent is compared with the answer in the provided knowledge document in the case of this study. QA metrics like the QA-Ref one have been analysed in evaluating performances of conversational agents as seen in Reddy et al. 2019 work [14], where QA performance was measured against selected reference answers.
  • QA-Cand: In scenarios where there could be multiple plausible answers from the reference document, QA-Cand metrics offer a good evaluation standard. It portrays a high tolerance for flexibility in the candidate responses which can be sourced from the custom documents in our case. Candidate-based evaluation has been explored in some open-domain QA tasks, for example in Kwiatkowski et al. 2019 work [7], where various inclusive answers were considered for evaluating their model performance.
For Fluency and Coherence:
  • BERT-Precision, BERT-Recall, and BERT-F1: These three rely on BERT-embeddings to capture semantic similarities between the chatbot response and the custom document for the case of BERT-Precision, the breadth and coverage of the response based on the custom document for BERT-Recall, and BERT-F1 which gives a balanced tradeoff between the precision (conciseness) and the recall (completeness) offering a more aligned semantic balance.
  • METEOR: This evaluates precision, recall, stemming, synonyms, and word alignment ensuring a linguistically diverse yet semantically accurate response, with a focus on fluency and flexibility as referenced with the custom document provided. In 2005, Banerjee et al. [2] proposed the METEOR as a more linguistically guided alternative to BLEU, developed to offer a better evaluation of language generation tasks.
  • Perplexity: This generally measures the effect of a probabilistic language model in the ability to predict the following words in a sequence scenario, by calculating the inverse of the probability of the set of test samples which are normalized by the frequency of the words. In 1992, Brown et al. [4] discussed perplexity as a key metric for evaluating n-gram models while in 2019, Radford et al. [13] proposed it as a metric for GPT models.
6 Experimental Results and Analysis
We first installed OLLAMA SDK and pulled the 5 models we needed for our work to our local computer. Our computer had a GPU of 12 GB RAM NVIDIA GEFORCE. After following the experimental set-up in section 5 above, we recorded the results from all the metrics used from the custom-designed web application including the response time after the query was passed to the model. We show the results in the two tables below with the first one consisting of the results without including the illocutionary force of the query, and the second table consisting of the results after the illocutionary force of the query was included.
Without Illocutionary Force added to User Query:
S/N Documents Model Bot Response BERT BERT BERT QA QA ROUGE-1 ROUGE-2 ROUGE-L METEOR Perplexity Response Time (sec) Precision Recall F1 Ref Cand 1 001 llama2:13b First 74.64 0.72 0.72 0.72 0.49 0.24 0.24 0.13 0.19 0.42 56.75 2 001 Tinyllama:latest First 26.55 0.68 0.68 0.68 0.49 0.39 0.18 0.07 0.12 0.35 26.69 3 001 Llama3-chatqa-latest First 74.33 0.84 0.84 0.84 0.49 0.37 0.43 0.21 0.38 0.49 54.7 4 001 Llama3-latest First 105.02 0.74 0.74 0.74 0.49 0.63 0.16 0.01 0.09 0.22 49.16 5 001 Mistral:latest First 109.38 0.57 0.57 0.57 0.49 0.55 0.19 0.04 0.12 0.30 28.64 6 001 llama2:13b Second 164.58 0.92 0.92 0.92 0.04 0.17 0.39 0.13 0.24 0.46 70.62 7 001 Tinyllama:latest Second 18.04 0.81 0.81 0.81 0.04 0.09 0.33 0.09 0.18 0.41 48.82 8 001 Llama3-chatqa-latest Second 62.80 0.90 0.90 0.90 0.04 0.31 0.25 0.08 0.18 0.26 29.12 9 001 Llama3-latest Second 85.69 0.86 0.86 0.86 0.04 0.17 0.19 0.04 0.13 0.28 55.14 10 001 Mistral:latest Second 130.64 0.79 0.79 0.79 0.04 -0.03 0.17 0.06 0.10 0.31 28.96 11 004 llama2:13b First 543.83 0.78 0.78 0.78 0.77 0.63 0.12 0.05 0.10 0.21 41.60 12 004 Tinyllama:latest First 57.02 0.84 0.84 0.84 0.77 0.03 0.10 0.00 0.08 0.15 1849.46 13 004 Llama3-chatqa-latest First - - - - - - - - - - - 14 004 Llama3-latest First 252.28 0.74 0.74 0.74 0.77 0.68 0.11 0.04 0.08 0.19 26.80 15 004 Mistral:latest First 279.70 0.80 0.80 0.80 0.77 0.65 0.14 0.06 0.11 0.22 29.15 16 004 llama2:13b Second 532.95 0.89 0.89 0.89 0.02 -0.01 0.15 0.09 0.11 0.40 45.45 17 004 Tinyllama:latest Second 52.84 0.93 0.93 0.93 0.02 0.03 0.19 0.05 0.19 0.20 2612.20 18 004 Llama3-chatqa-latest Second - - - - - - - - - - - 19 004 Llama3-latest Second 237.64 0.89 0.89 0.89 0.02 0.10 0.25 0.05 0.16 0.29 35.32 20 004 Mistral:latest Second 298.41 0.82 0.82 0.82 0.02 0.07 0.15 0.06 0.10 0.23 34.43
With Illocutionary Force added to User Query:
S/N Documents Model Bot Response BERT BERT BERT QA QA ROUGE-1 ROUGE-2 ROUGE-L METEOR Perplexity Response Time (sec) Precision Recall F1 Ref Cand 1 001 llama2:13b First 212.15 0.79 0.79 0.79 0.49 0.51 0.22 0.07 0.12 0.33 72.52 2 001 Tinyllama:latest First 24 0.74 0.74 0.74 0.49 0.48 0.23 0.10 0.13 0.41 32.90 3 001 Llama3-chatqa-latest First 70.92 0.64 0.64 0.64 0.49 0.68 0.24 0.02 0.14 0.24 58.74 4 001 Llama3-latest First 92.45 0.66 0.66 0.66 0.49 0.63 0.19 0.02 0.11 0.28 60.24 5 001 Mistral:latest First 107.49 0.73 0.73 0.73 0.49 0.45 0.24 0.07 0.13 0.37 28.67 6 001 llama2:13b Second 244.50 0.79 0.79 0.79 0.043 0.066 0.22 0.07 0.12 0.33 73.07 7 001 Tinyllama:latest Second 28.60 0.69 0.69 0.69 0.04 0.03 0.17 0.07 0.10 0.29 29.29 8 001 Llama3-chatqa-latest Second 64.05 0.98 0.98 0.98 0.07 0.60 0.14 0.0 0.14 0.10 70.82 9 001 Llama3-latest Second 80.86 0.87 0.87 0.87 0.04 0.11 0.23 0.03 0.15 0.32 20.44 10 001 Mistral:latest Second 108.78 0.83 0.83 0.83 0.04 0.09 0.25 0.06 0.17 0.37 32.75 11 004 llama2:13b First 611.63 0.83 0.83 0.83 0.77 0.69 0.13 0.05 0.09 0.25 39.09 12 004 Tinyllama:latest First 59.09 0.79 0.79 0.79 0.77 0.15 0.13 0.03 0.09 0.16 2552.25 13 004 Llama3-chatqa-latest First - - - - - - - - - - - 14 004 Llama3-latest First 350.84 0.80 0.80 0.80 0.77 0.64 0.11 0.03 0.07 0.19 31.40 15 004 Mistral:latest First 375.14 0.76 0.76 0.76 0.77 0.61 0.12 0.04 0.09 0.28 51.88 16 004 llama2:13b Second 613.46 0.82 0.82 0.82 0.02 -0.067 0.15 0.05 0.08 0.22 51.75 17 004 Tinyllama:latest Second 67.30 0.86 0.86 0.86 0.77 0.03 0.09 0.05 0.06 0.25 583.08 18 004 Llama3-chatqa-latest Second - - - - - - - - - - - 19 004 Llama3-latest Second 330.94 0.86 0.86 0.86 0.019 0.017 0.21 0.042 0.12 0.28 104.75 20 004 Mistral:latest Second 329.08 0.83 0.83 0.83 0.02 -0.019 0.20 0.07 0.12 0.32 41.03
6.1 Result Analysis
In the analysis below:
F $\Rightarrow$ indicates Faster
1 $\Rightarrow$ indicates a better performance
0 $\Rightarrow$ indicates a poor performance
S $\Rightarrow$ indicates a similar performance
Analysis of the two results based on first bot responses in Document 001:
Without Illocutionary Force With Illocutionary Force
llama2 Tinyllama Llama3-chatqa Llama3 Mistral llama2 Tinyllama Llama3-chatqa Llama3 Mistral
:13b :latest -latest -latest :latest :13b :latest -latest -latest :latest
Response Time F - - - - - F F F F
BERT-Precision 0 0 1 1 0 1 1 0 0 1
BERT-Recall 0 0 1 1 0 1 1 0 0 1
BERT-F1 0 0 1 1 0 1 1 0 0 1
QA-Ref S S S S S S S S S S
QA-Cand 0 0 0 S 1 1 1 1 S 0
ROUGE-1 1 0 1 0 0 0 1 0 1 1
ROUGE-2 1 0 1 0 0 0 1 0 1 1
ROUGE-L 1 0 1 0 0 0 1 0 1 1
METEOR 1 0 1 0 0 0 1 0 1 1
Perplexity 1 1 1 1 1 0 0 0 0 0
Avg Total 5 1 8 4 2 4 8 1 4 7
Table 4: Performance metrics for various models for the bot’s first response using the first document
From the table above, we analysed the following:
  1. 1.
    Response Time: It is observed that the response time of the chat agent was faster in all the models for the experiment which included the illocutionary force except for the llama2:13b model in which it was faster in the experiment without the illocutionary force included in the user query.
  2. 2.
    BERT-Precision: It is observed that the semantic similarities based on the BERT-embeddings were better in Llama3-chatqa-latest and Llama3-latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
  3. 3.
    BERT-Recall: It is observed that the breadth and coverage similarities based on the BERT-embeddings of the chat agent response and the custom document were better in Llama3-chatqa-latest and Llama3-latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
  4. 4.
    BERT-F1: It is observed that a balanced trade-off between the precision and recall which gives a better semantic balance based on the BERT-embeddings of the chat agent response and the custom document were better in Llama3-chatqa-latest and Llama3-latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
  5. 5.
    QA-Ref: Comparing the chat agent response typically with answers in the knowledge document, it is observed that all the models in both the experiment without and with the illocutionary force included had the same performance indicating a balanced metric.
  6. 6.
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that llama2:13b, Tinyllama:latest, Llama3-chatqa-latest performed better in the experiment with the illocutionary force included, performed same with llama3-latest, and performed better in mistral:latest for the experiment without the illocutionary force.
  7. 7.
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better in llama2:13b and Llama3-chatqa-lates models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
  8. 8.
    ROUGE-2: It is observed that elements of Bigram indicating coherence in the chat agent response based on the custom document were better in llama2:13b and Llama3-chatqa-lates models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
  9. 9.
    ROUGE-L: It is observed that the longest common subsequence between the chat agent response and the custom document was better in llama2:13b and Llama3-chatqa-latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
  10. 10.
    METEOR: It is observed that a better linguistically diverse yet semantically response from the chat agent with respect to the custom document, was better in llama2:13b and Llama3-chatqa-latest models for the experiment conducted without including the illocutionary force, and better in the other models for the experiment conducted with the inclusion of the illocutionary force with the user query.
  11. 11.
    Perplexity: The ability to predict the following words in a sequence scenario was better in all the models for the experiment conducted without including the illocutionary force with a lower perplexity value as compared to same models for the experiment conducted with the inclusion of the illocutionary force with the user query.
  12. 12.
    Avg Total: Taking a record of the performance of the models across all metrics and excluding both the response time and the results of the similar values, it is observed that: the llama2:13b model performs better with an average score of 5 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a score of 4; the Tinyllama:latest model performs better with an average score of 8 in the experiment with the illocutionary force against same model for the experiment without the illocutionary force which had a score of 1; the Llama3-chatqa-latest model performs better with an average score of 8 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a score of 1; the Llama3-latest model had similar score of 4 in the experiment with the illocutionary force and for the experiment without the illocutionary force which also had a score of 4; the Mistral:latest model performs better with an average score of 7 in the experiment with the illocutionary force against same model for the experiment without the illocutionary force which had a score of 2.
Analysis of the two results based on second bot responses in Document 001:
Without Illocutionary Force With Illocutionary Force
llama2 Tinyllama Llama3-chatqa Llama3 Mistral llama2 Tinyllama Llama3-chatqa Llama3 Mistral
:13b :latest -latest -latest :latest :13b :latest -latest -latest :latest
Response Time F F F - - - - - F F
BERT-Precision 1 1 0 0 0 0 0 1 1 1
BERT-Recall 1 1 0 0 0 0 0 1 1 1
BERT-F1 1 1 0 0 0 0 0 1 1 1
QA-Ref 0 S 0 S S 1 S 1 S S
QA-Cand 1 1 0 1 0 0 0 1 0 1
ROUGE-1 1 1 1 0 0 0 0 0 1 1
ROUGE-2 1 1 1 1 S 0 0 0 0 S
ROUGE-L 1 1 1 0 0 0 0 0 1 1
METEOR 1 1 1 0 0 0 0 0 1 1
Perplexity 1 0 1 0 1 0 1 0 1 0
Avg Total 9 8 5 2 1 1 1 5 7 7
Table 5: Performance metrics for various models for the bot’s second response using the first document
From the table above, we analysed the following:
  1. 1.
    Response Time: It is observed that the response time of the chat agent was faster in Llama3-latest and Mistral:latest for the experiment which included the illocutionary force and was faster for the other models in the experiment without the illocutionary force included in the user query.
  2. 2.
    BERT-Precision: It is observed that the semantic similarities based on the BERT-embeddings were better in llama2:13b and Tinyllama:latest models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
  3. 3.
    BERT-Recall: It is observed that the breadth and coverage similarities based on the BERT-embeddings of the chat agent response and the custom document were better in llama2:13b and Tinyllama:latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
  4. 4.
    BERT-F1: It is observed that a balanced trade-off between the precision and recall which gives a better semantic balance based on the BERT-embeddings of the chat agent response and the custom document were better in llama2:13b and Tinyllama:latest models in the experiment without including the illocutionary force, and was better in the rest of the other models in the experiment which included the illocutionary force.
  5. 5.
    QA-Ref: Comparing the chat agent response typically with answers in the knowledge document, it is observed that Tinyllama:latest, Llama3-latest, and Mistral:latest models in both the experiment without and with the illocutionary force included had the same performance indicating a balanced metric, but was better llama2:13b, and Llama3-chatqa-latest in the experiment with the inclusion of the illocutionary force as compared to the experiment without the illocutionary force.
  6. 6.
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that Llama3-chatqa-latest, and Mistral:latest performed better in the experiment with the illocutionary force included, and performed better in other models for the experiment without the illocutionary force.
  7. 7.
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better in llama2:13b, Tinyllama:latest, and Llama3-chatqa-latest models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
  8. 8.
    ROUGE-2: It is observed that elements of bigram indicating coherence in the chat agent response based on the custom document were better in all the models in the experiment without including the illocutionary force, and was same in only the Mistral:latest model in the experiment which included the illocutionary force.
  9. 9.
    ROUGE-L: It is observed that the longest common subsequence between the chat agent response and on the custom document was better in llama2:13b, Tinyllama:latest, and Llama3-chatqa-latest models in the experiment without including the illocutionary force, and were better in the rest of the other models in the experiment which included the illocutionary force.
  10. 10.
    METEOR: It is observed the a better linguistically diverse but yet sementically response from the chat agent with respect to the custom document, was better in llama2:13b, Tinyllama:latest, and Llama3-chatqa-latest models for the experiment conducted without including the illocutionary force, and better in the other models for the experiment conducted with the inclusion of the illocutionary force with the user query.
  11. 11.
    Perplexity: The ability to predict the following words in a sequence scenario was better in llama2:13b, Llama3-chatqa-latest, and Mistral:latest models for the experiment conducted without including the illocutionary force with a lower perplexity value and better in other models for the experiment conducted with the inclusion of the illocutionary force with the user query.
  12. 12.
    Avg Total: Taking a record of the performance of the models across all metrics and excluding both the response time and the results of the similar values, it is observed that: the llama2:13b model performs better with an average score of 9 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a score of 1; the Tinyllama:latest model performs better with an average score of 8 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a score of 1; the Llama3-chatqa-latest model performs same with an average score of 5 in the experiment without the illocutionary force against same model for the experiment with the illocutionary force which had a similar score of 5; the Llama3-latest model performed better with a score of 7 in the experiment with the illocutionary force and for the experiment without the illocutionary force with a score of 2; the Mistral:latest model performs better with an average score of 7 in the experiment with the illocutionary force against same model for the experiment without the illocutionary force which had a score of 1.
Analysis of the two results based on first bot responses in Document 004
Without Illocutionary Force With Illocutionary Force
llama2 Tinyllama Llama3-chatqa Llama3 Mistral llama2 Tinyllama Llama3-chatqa Llama3 Mistral
:13b :latest -latest -latest :latest :13b :latest -latest -latest :latest
Response Time F F - F F - - - - -
BERT-Precision 0 1 - 0 1 1 0 - 1 0
BERT-Recall 0 1 - 0 1 1 0 - 1 0
BERT-F1 0 1 - 0 1 1 0 - 1 0
QA-Ref S S - S S S S - S S
QA-Cand 0 0 - 1 1 1 1 - 0 0
ROUGE-1 0 0 - S 1 1 1 - S 0
ROUGE-2 S 0 - 1 1 S 1 - 0 0
ROUGE-L 1 0 - 1 1 0 1 - 0 0
METEOR 0 0 - S 0 1 1 - S 1
Perplexity 0 1 - 1 1 1 0 - 0 0
Avg Total 1 4 - 4 8 7 5 - 3 1
Table 6: Performance metrics for various models for the bot first response using the second document
From the table above, we analysed the following:
  1. 1.
    Response Time: It is observed that the response time for Llama2:13b, Tinyllama:latest Llama3-latest, and Mistral:latest is faster in the experiment conducted without the illocutionary force as compared with the same models in the experiment with the illocutionary force.
  2. 2.
    BERT-Precision: It is observed that the precision metric based on BERT-embeddings showed a better score for Tinyllama:latest and Mistral:latest in the experiment conducted without including the illocutionary force and a better score was observed in the Llama3-latest, and Llama2:13b models for the experiment that included the illocutionary force in the user query.
  3. 3.
    BERT-Recall: It is observed that the recall metric based on BERT-embeddings is similar to the precision results with better scores noted for Tinyllama:latest and Mistral:latest in the experiment without including the illocutionary force, and better scores for Llama3-latest and Llama2:13b in the experiment with the illocutionary force included with the user query.
  4. 4.
    BERT-F1: It is observed that the balanced precision and recall showed similar results to BERT-Recall and Precision, indicating a better score for Tinyllama:latest and Mistral:latest in the experiment without illocutionary force, and better scores for Llama3-latest and Llama2:13b with the illocutionary force included with the user query.
  5. 5.
    QA-Ref: Comparing the chat agent response typically with answers in the knowledge document, it is observed that the performance was similar across all models in both experiments, indicating a high similarity in this metric.
  6. 6.
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that Llama3-latest and Mistral:latest showed better results in the experiment without the illocutionary force, and better results for Llama2:13b and Tinyllama:latest in the experiment when the illocutionary force was included with the user query.
  7. 7.
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better for Mistral:latest in the experiment without the illocutionary force, while Llama2:13b and Tinyllama:latest models showed better scores in the experiment with the illocutionary force included, and a similar score for llama3-latest in both experiments.
  8. 8.
    ROUGE-2: It is observed that bigram coherence was better for Llama3-latest and Mistral:latest in the experiment without the illocutionary force, but was observed as similar for Llama2:13b across both experiments and better in Tinyllama:latest with the inclusion of the illocutionary force.
  9. 9.
    ROUGE-L: It is observed that the longest common subsequence metric was better for Llama2:13b, Llama3-latest, and Mistral:latest in the experiment conducted without the illocutionary force. However, Tinyllama:latest was the only model to achieve a better result in the experiment with the illocutionary force included in the user query.
  10. 10.
    METEOR: It is observed that linguistic diversity and semantic accuracy in responses showed a better score for Llama2:13b, Tinyllama:latest, and Mistral:latest in the experiment with the illocutionary force, but was observed to be similar for Llama3-latest in both experiments.
  11. 11.
    Perplexity: It is observed that the predictive ability of the models indicated better values for Tinyllama:latest, Llama3-latest and Mistral:latest in the experiment without the illocutionary force, while the same metric achieved a better score in Llama2:13b for the experiment with the illocutionary force.
  12. 12.
    Avg Total: Taking a record of the performance of the models across all metrics, it was observed that the Mistral:latest model performed best with a total score of 8 in the experiment without the illocutionary force and a much lower score of 1 in the experiment with the illocutionary force. Similarly, Tinyllama:latest obtained a score of 4 in the experiment without the illocutionary force but increased to 5 when the illocutionary force was included. Llama2:13b had a significant difference in performance, with a score of 1 in the experiment without illocutionary force but improved to 7 when it was included in the user query. Finally, Llama3-latest showed moderate performance, scoring 4 without the illocutionary force and 3 when it was included in the user query.
Analysis of the two results based on second bot responses in Document 004
Without Illocutionary Force With Illocutionary Force
llama2 Tinyllama Llama3-chatqa Llama3 Mistral llama2 Tinyllama Llama3-chatqa Llama3 Mistral
:13b :latest -latest -latest :latest :13b :latest -latest -latest :latest
Response Time F F - F F - - - - -
BERT-Precision 1 1 - 1 0 0 0 - 0 1
BERT-Recall 1 1 - 1 0 0 0 - 0 1
BERT-F1 1 1 - 1 0 0 0 - 0 1
QA-Ref S 0 - 1 S S 1 - 0 S
QA-Cand 1 S - 1 1 0 S - 0 0
ROUGE-1 S 1 - 1 0 S 0 - 0 1
ROUGE-2 1 S - 1 0 0 S - 0 1
ROUGE-L 1 1 - 1 0 0 0 - 0 1
METEOR 1 0 - 1 0 0 1 - 0 1
Perplexity 1 0 - 1 1 0 1 - 0 0
Avg Total 8 5 - 10 2 0 3 - 0 7
Table 7: Performance metrics for various models for the bot’s second response using the second document
From the table above, we analysed the following:
  1. 1.
    Response Time: It is observed that the response time for Llama2:13b, Tinyllama:latest Llama3-latest, and Mistral:latest is faster in the experiment conducted without the illocutionary force as compared with the same models in the experiment with the illocutionary force.
  2. 2.
    BERT-Precision: It is observed that the precision metric based on BERT-embeddings showed a better score for llama2:13b, Tinyllama:latest and Llama3-latest in the experiment conducted without including the illocutionary force and a better score was observed in Mistral:latest model for the experiment that included the illocutionary force in the user query.
  3. 3.
    BERT-Recall: It is observed that the recall metric based on BERT-embeddings is similar to the precision results with better scores for llama2:13b, Tinyllama:latest and Llama3-latest in the experiment conducted without including the illocutionary force and a better score was observed in Mistral:latest model for the experiment that included the illocutionary force in the user query.
  4. 4.
    BERT-F1: It is observed that the balanced precision and recall showed similar results to BERT-Recall and Precision, indicating better scores for llama2:13b, Tinyllama:latest and Llama3-latest in the experiment conducted without including the illocutionary force and a better score was observed in Mistral:latest model for the experiment that included the illocutionary force in the user query.
  5. 5.
    QA-Ref: Comparing the chat agent response typically with answers in the knowledge document, it is observed that the performance was similar for llama2:13b and Mistral:latest in both experiments, indicating a high similarity in this metric but Llama3-latest was better in the first experiment, and Tinyllama:latest better in the second experiment.
  6. 6.
    QA-Cand: When we observed the flexibility of the response with respect to the custom document provided, it was seen that Llama2:13b, Llama3-latest and Mistral:latest showed better results in the experiment without the illocutionary force, and the same result for Tinyllama:latest in both experiments.
  7. 7.
    ROUGE-1: It is observed that elements of unigram keywords in the chat agent response found in the custom document were better for Tinyllama:latest, and llama3-latest in the experiment without the illocutionary force, while Mistral:latest model showed better scores in the experiment with the illocutionary force included, and a similar score for Llama2:13b in both experiments.
  8. 8.
    ROUGE-2: It is observed that bigram coherence was better for Llama2:13b and Llama3-latest in the experiment without the illocutionary force, but was observed as similar for Tinyllama:latest across both experiments and better in Mistral:latest with the inclusion of the illocutionary force.
  9. 9.
    ROUGE-L: It is observed that the longest common subsequence metric was better for Llama2:13b, Llama3-latest, and Tinyllama:lates in the experiment conducted without the illocutionary force. However, Mistral:latest was the only model to achieve a better result in the experiment with the illocutionary force included in the user query.
  10. 10.
    METEOR: It is observed that linguistic diversity and semantic accuracy in responses showed a better score for Llama2:13b, and Llama3-latest in the experiment with the illocutionary force, but was observed to be better for Tinyllama:latest, and Mistral:latest in the experiment with the illocutionary force included in the user query.
  11. 11.
    Perplexity: It is observed that the predictive ability of the models indicated better values for Llama2:13b, Llama3-latest and Mistral:latest in the experiment without the illocutionary force, while the same metric achieved a better score in Tinyllama:latest for the experiment with the illocutionary force.
  12. 12.
    Avg Total: Taking a record of the performance of the models across all metrics, it was observed that the Llama3-latest model performed best with a total score of 10 in the experiment without the illocutionary force and a much lower score of 0 in the experiment with the illocutionary force. Similarly, llama2:13b obtained a score of 8 in the experiment without the illocutionary force but dropped to 0 when the illocutionary force was included. Tinyllama:latest showed a difference in performance, with a score of 5 in the experiment without illocutionary force but dropped to 3 when it was included in the user query. Finally, Mistral:latest showed a performance of 2 without the illocutionary force and was improved to 7 when it was included in the user query.
7 Conclusion and Future work
The results from this research unlock several potentials for the inclusion of communicative intent (illocutionary forces) which have been overlooked by computational linguistics in improving the responses given by conversational agents in an HCI setting. The illocutionary forces explored in this research originate from a dialogue domain making it very relevant as an additional linguistic feature layer which is not currently embedded in the context of transformer models. To improve on this work, there is a need to explore larger encoder-decoder models with larger parameters and explore the impact of repeating these experiments on a more robust GPU.